OpenStudy (anonymous):

f f(x) = 4x – 6 and g(x) = 2x, find f(x)*g(x)

OpenStudy (anonymous):

*if

OpenStudy (anonymous):

(4x-6)(2x)= 8x^2-12x

OpenStudy (anonymous):

thank you ! Can you help with more ?

OpenStudy (anonymous):

sure! just post them here

OpenStudy (anonymous):

If f(x)=2x^2-4x and g(x) = 5 – 2x, evaluate f(x) – g(x) for x = 5.

OpenStudy (anonymous):

You got it?

OpenStudy (anonymous):

f(x)=2x^2-4x, for x=5 f(x))=50-20 f(x)=30 g(x)=5-2x g(x)=5-10 g(x)=-5 f(x)-g(x)= 30 - (-5) f(x)-g(x)= 35

OpenStudy (anonymous):

You dont have to type it all. just the answer :)

OpenStudy (anonymous):

oh lol. oh well

OpenStudy (anonymous):

But thank you soo much .. lol hate summer school ... If f(x)=-2x^2+1 and g(x) = 3x, evaluate f(x)/g(x) for x = 2.

OpenStudy (anonymous):

- 7/6

OpenStudy (anonymous):

The revenue of a company is represented by r(x) and the expenses of the company are represented by e(x). What would you do to find the profit, p(x)?

OpenStudy (anonymous):

revenue - expenses = profit

OpenStudy (anonymous):

this is the whole question

OpenStudy (anonymous):

The revenue of a company is represented by r(x) and the expenses of the company are represented by e(x). What would you do to find the profit, p(x)? A. r(x) + e(x) = p(x) B. r(x) – e(x) = p(x) C. e(x) – r(x) = p(x) D. r(x) / e(x) = p(x)

OpenStudy (anonymous):

haha its B

OpenStudy (anonymous):

If f(x) = 4x – 7 and g(x) = 2x + 4, evaluate f(x) + g(x) for x = –3.

OpenStudy (anonymous):

-21

OpenStudy (anonymous):

If f(x) = 4x – 5 and g(x) = 3 – 2x, find f(x) + g(x)

OpenStudy (anonymous):

2x-2

OpenStudy (anonymous):

^^^

OpenStudy (anonymous):

one sec lol

OpenStudy (anonymous):

haha okay

OpenStudy (anonymous):

i think its (x^2 - 3) / (3x) but that could be wrong

OpenStudy (anonymous):

I gotta go though, I have lacrosse practice in 20 minutes haha gotta get ready. hope that helped!

OpenStudy (anonymous):

Aw, thank you so much though!!